OpenStudy (anonymous):

how can you solve 4x+6w=1000

OpenStudy (xxdragoonxx):

6w+4x-6w=1000+-6w 4x=-6w+1000 4x/4=-6w+1000/4 x=-3/2w+250 6w+4x+-4x=1000+4-4x 6w=-4x+1000 6w/6=-4x+1000/6 w=-2/3x+500/3

TheSmartOne (thesmartone):

\(\Huge\cal\color{blue}{Welcome~to~Open~Study!!}\) It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

What is your question asking for us to solve for???